# Learning Top-$k$ Subtask Planning Tree based on Discriminative Representation Pre-training for Decision Making

### v1

@@ 4.2. Results on Subtask Representaion
Figure 8. The training curve comparisons on six complicated tasks, averaged over 10 independent running seeds.
4.3. Results on Training Curves
To evaluate our method in complicated and dynamic environments, we conduct various experiments on multiple difficulty levels of BabyAI, elaborated in Appendix A.3. These levels require multiple competencies to solve them. For example, GotoSeq only requires navigating a room or maze, recognizing the location, and understanding the sequence subtasks while BossLevel requires mastering all competencies including navigating, opening, picking up, putting next, and so on.
Here, we report the average episode reward by training steps for each level, averaged over 10 independent running seeds to plot the reward curve with standard deviation. In GoToSeqS5R2 which is the relatively simplest level in the six, all the test rewards keep increasing as the training steps increase and almost reach the best finally. In other complicated tasks, our algorithm consistently obtains higher rewards than all the baselines in the different levels of BabyAI shown in Figure 8, while the popular SOTA like PPO and SAC almost fail. This indicates that our method is capable of macro-planning and robust enough to handle the complicated tasks by decomposing the overall process into a sequence of execution subtasks.
Although we only trained 10M timesteps due to the limitation of computing resources, our method has a higher starting point than the baseline model, and it has been rising. However, the gain of the baselines at the same timestep is not obvious in most cases. This significant improvement is mainly due to two aspects. On the one hand, the adequate knowledge representation facilitates the inference capacity of our model and gives a high-level starting point for warming up the training. On the other hand, the learned subtask planning strategy can make rational decisions while confronting various complicated situations.
4.4. Results on Execution Performance
@@
Table 1. Execution performance on various complicated tasks. We report the average reward with standard deviations of 100 testing episodes with best values in bold.
Ours CORRO PPO SAC SGT
GoToSeq
0.30$\pm$0.07
0.23$\pm$ 0.06
0.14$\pm$ 0.04
0.05$\pm$ 0.04
0.15$\pm$ 0.08
GoToSeqS5R2
0.91$\pm$ 0.02
0.84$\pm$ 0.02
0.84$\pm$ 0.03
0.85$\pm$ 0.02
0.91$\pm$ 0.04
SynthSeq
0.22$\pm$ 0.05
0.16$\pm$ 0.05
0.12$\pm$ 0.05
0.12$\pm$ 0.04
0.18$\pm$ 0.05
BossLevel
0.26$\pm$ 0.05
0.15$\pm$ 0.04
0.14$\pm$ 0.05
0.12$\pm$ 0.04
0.19$\pm$ 0.06
BossLevelNoUnlock
0.24$\pm$ 0.04
0.17$\pm$ 0.04
0.10$\pm$ 0.05
0.04$\pm$ 0.03
0.11$\pm$ 0.05
MiniBossLevel
0.29$\pm$ 0.06
0.26$\pm$ 0.05
0.25$\pm$ 0.05
0.06$\pm$ 0.03
0.14$\pm$ 0.06
4.5. Results on Various Tree Width
After the verification of the training and execution performance of our technique, we perform in-depth ablation studies regarding the width of the planning tree shown in Figure 9. Here, we freeze the tree depth as $3$, same as the main experiments. Additional experimental results about other scenarios can be found in Appendix A.6.
@@ 4.5. Results on Various Tree Width
(b) GotoSeq
Figure 9. The ablations on various tree width.
The horizontal axis represents the training steps, and the vertical axis is the average episode reward with a standard deviation, averaged across 10 independent running seeds. According to our intuition, the wider the tree is, the more diverse the subtask distribution is. However, according to the experimental results, we find that the gain is not the largest when the width of the tree is the largest. We speculate that as the width of the tree is larger, more information can be selected, but the pertinence of problem-solving is not strong, which is easy to cause confusion. Therefore, it is critical to choose an appropriate tree width and it is our future direction to conduct the adaptive selection.
4.6. Results on Various Tree Depth
As shown in Figure 10, we further provide in-depth ablation studies to understand the behavior of our planning tree depending on various tree depths. Here, we freeze the tree width as $2$, the same as the main experiments. Then we provide more ablation studies in Appendix A.7. The depth represents the capacity for predicting the future of our planning tree. As shown in Figure 10, the results show that when the depth $N=2$ or $N=3$, our model wins the comparable performance also. However, the curves about $N=4$ shows the degraded performance than others. We analyze that the prediction error about the future at every step keeps accumulating, so it may make the model go in a bad direction although the foresight ability can help the model make a global decision.
@@ 4.6. Results on Various Tree Depth
(b) GotoSeq
Figure 10. The ablations on various tree depth.
4.7. Results on Subtask Execution Statistics
As shown in Figure 11, we visualize the subtask proportion on six complicated tasks after testing 1000 episodes respectively. It is surprisingly consistent with the pre-defined task missions. It basically corresponds to predefined tasks, which intuitively justifies the rationality of our approach. For example, GoToSeq and GoToSeqS5R2 only require the competency of executing some sequences about“Go To”, our model focuses on this subtask during the execution in most cases. There exists some error that other tasks will be selected occasionally, and we speculate that it is because of the inaccuracy of forward prediction during expanding the depth of the tree. Our model executes subtasks at roughly the same frequency when dealing with the other four tasks which require comprehensive competencies.
Figure 11. The bars represent the mean proportion of each subtask executed by our trained model over 1000 episodes.
5. Conclusion
In this paper, we have considered the difficulties of task planning to solve complicated problems, and propose a two-stage planning framework. The multiple-encoder and individual-predictor regime can extract discriminative subtask representations with dynamic knowledge from priors. The top-k subtask planning tree, which customizes subtask execution plans to guide decisions, is important in complicated tasks. Generalizing our model to unseen and complicated tasks demonstrates significant performance and exhibits greater potential to deploy it to real-life complicated tasks. More ablation studies also demonstrate the capacity of our planning tree.
For future works, it would be interesting to automatically learn more explainable subtasks and make influential and global planning as a reference to human decisions in real-life applications.
Appendix A Appendix
A.1. Detailed Statistics about the Collected Data
The data collection statistics are shown as Table 2, of which 1/3 is collected from the converged PPO policy, 1/3 is collected from the intermediate trained checkpoint of PPO, and the last 1/3 is from the random policy.
Subtask #Total Episodes #Total Transitions
PickupLoc 3000 210375
PutNext 3000 559201
Goto 3000 218191
Open 3000 149349
Table 2. The data collection statistics
A.2. Detailed Description about the Intuition for d-UCB
Figure 12. Two instance trees are visualized to verify the effect of d-UCB described in Section 3.4. To ignore the effect of UCB value, we set the value of all nodes as $1$.
@@ A.2. Detailed Description about the Intuition for d-UCB
As shown in Figure 12(a), MS will select path $1$, which obviously places a greater weight on the future with greater uncertainty. However, our d-UCB will choose path $2$, which tends to maximize the immediate benefit by discounting the future uncertain score, which is consistent with our intuition.
Moreover, as shown in Fig. 12(b), not like MS tending to pick the edge with the highest probability, our d-UCB will choose path $3$. Our d-UCB also considers the optimal subtask to perform in the future when the nearest paths exhibit similar transition probabilities.
A.3. Detailed Description about Complicated Tasks
(a) ‘BabyAI-BossLevel-v0’
(b) ‘BabyAI-BossLevelNoUnlock-v0’
(c) ‘BabyAI-MiniBossLevel-v0’
(d) ‘BabyAI-GoToSeq-v0’
(e) ‘BabyAI-GoToSeqS5R2-v0’
(f) ‘BabyAI-SynthSeq-v0’
Figure 13. The diagrams for six complicated tasks in BabyAI.
BossLevel
The level requires the agent to implement any union of instructions produced according to Baby Language grammar, it involves the content of all other levels. For example, “open a door and pick up the green key after you go to a ball and put a red box next to the red ball” is a command of the level. This level requires all the competencies including navigating, unblocking, unlocking, opening, going to, picking up, putting next, and so on.
BossLevelNoUnlock
The doors of all rooms are unlocked which the agent can go through without unlocking by a corresponding key. For example, the command “open a purple door, then go to the yellow box and put a yellow key next to a box” doesn’t need the agent to get a key before opening a door. This level requires all the competencies like BossLevel except unlocking.
MiniBossLevel
The level is a mini-environment of BossLevel with fewer rooms and a smaller size. This level requires all the competencies like BossLevel except unlocking.
GoToSeq
The agent needs to complete a series of go-to-object commands in sequence. For example, the command ”go to a grey key and go to a purple key, then go to the green key” only includes go-to instructions. GotoSeq only requires navigating a room or maze, recognizing the location, and understanding the sequence subtasks.
GoToSeqS5R2
The missions are just like GoToSeq, which requires the only competency of “go to”. Here, S5 and R2 refer to the size and the rows of room, respectively.
SynthSeq
The commands, which combine different kinds of instructions, must be executed in order.For example, ”open the purple door and pick up the blue ball, then put the grey box next to a ball”. This level requires the competencies like GotoSeq except inferring whether to unlock or not.
A.4. Network Architecture
There is a summary of all the neural networks used in our framework about the network structure, layers, and activation functions.
@@ A.4. Network Architecture
Please note that ’-’ denote that we refer readers to check the open source repository about FiLMedBlock and CausalSelfAttention, see https://github.com/caffeinism/FiLM-pytorch and https://github.com/sachiel321/Efficient-Spatio-Temporal-Transformer, respectively.
A.5. Parameter Settings
There are our hyper-parameter settings for the training of distinguishable subtask representation and the tree-based policy, shown in Table A.5 and Table A.5, respectively.
Description Value
lr $3*10^{-4}$
contrastive batch size $64$
negative pairs number $16$
task embedding size $5$
$\lambda_{r}$ $0.9$
$\lambda_{s}$ $1.0$
encoder hidden size $[256,256,128,128,32]$
decoder hidden size $[64,128]$
Table 4. Task Representation Hyper-paramters
Description Value optimizer Adam $\alpha$ $10^{-4}$ $\beta_{1}$ $0.9$ $\beta_{2}$ $0.999$ $\varepsilon$ $10^{-5}$ seed $[0,10)$ number of process $64$ total frames $10M$ lr $10^{-4}$ batch size $1280$ eval interval $1$ eval episodes $500$ image embedding $128$ memory of LSTM $128$ query hidden size $[32,16,8]$ number of head $1$ discounted factor of UCB $0.9$ $m$-step predictor $5$ Table 5. Tree-based Policy Hyper-paramters
(a) BossLevel
(b) BossLevelNoUnlock
(c) MiniBossLevel
(d) GoToSeq
(e) GoToSeqS5R2
(f) SynthSeq
Figure 14. The ablation study of tree width for other four complicated tasks in BabyAI.
A.6. Additional Experimental Results on Tree Width
As shown in Figure 14, the other four experiments basically presented results consistent with the analysis in the main text. Among them, the experimental environment of GoToSeqSR2 is relatively simpler, and the performance of our model basically reaches the optimal trend, so the ablation study of tree width does not seem to work. We believe that our model is capable of handling complex tasks and is not very discriminative for simple tasks. In addition, for MiniBossLevel, the experiment with a tree width of $2$ achieves the best performance. It indicates that our model can already show a good performance improvement in a slightly simpler environment with a tree width of $2$.
A.7. Additional Experimental Results on Tree Depth
As shown in Figure 15, we plot the training curves of the other four complicated tasks freezing the width as $2$. Combining the experiments in the main text, these results all show a common phenomenon: the larger the value of $N$, the worse the performance of our model. This cumulative effect of errors shows that the accuracy of our tree-building prediction model needs to be improved, and it is also a defect of our model. We will focus on how to improve the forward prediction ability of the model in future work. In addition, changing the depth of the tree in the scene GoToSeqS5R2 still cannot reflect the difference of the model. This shows that our model can achieve excellent performance even if our tree prediction module is not very accurate. On the one hand, this success may depend on the strong fault tolerance of our pre-trained representation model, and the generated representations can correctly guide model learning; on the other hand, it may depend on the design of our subtask termination function, which can stop the subtask in time when the task mismatch is found.
(a) BossLevel
(b) BossLevelNoUnlock
(c) MiniBossLevel
(d) GoToSeq
(e) GoToSeqS5R2
(f) SynthSeq
Figure 15. The ablation study of tree depth for other four complicated tasks in BabyAI.

### v2

@@ 5.2 Results on Subtask Representation
Figure 7: Training curve comparisons on six complicated tasks, averaged over 10 independent running seeds.
5.3 Results on Training Curves
We evaluate our method on multiple difficulty levels of BabyAI, which require mastering various competencies, including navigation, location recognition, sequence understanding, object manipulation, and more. For example, GotoSeq only requires navigating a room or maze, recognizing the location, and understanding the sequence subtasks while BossLevel requires mastering all competencies including navigating, opening, picking up, putting next, and so on. We conduct experiments on these levels to test the effectiveness of our algorithm in complicated and dynamic environments, as elaborated in Appendix B.
Here, we report the average episode reward by training steps for each level, averaged over 10 independent running seeds to plot the reward curve with standard deviation. In GoToSeqS5R2 which is the relatively simplest level of the six, all the test rewards keep increasing as the training steps increase and finally almost reach the best. It is worth noting that although LEAP had a relatively good starting point in its early stages, the upwards trend was not evident in the later stages. We speculate that this is because LEAP, during its training, had encountered sufficient historical data, giving the model a higher starting point. However, the model’s capacity in the later stages seemed relatively weaker, preventing it from achieving even higher performance. In other complicated tasks, our algorithm consistently obtains higher rewards than all the baselines in the different levels of BabyAI shown in Fig. 7, while the popular SOTA algorithms, such as PPO and SAC almost fail. This indicates that our method is capable of macroplanning and robust enough to handle complicated tasks by decomposing the overall process into a sequence of execution subtasks.
Although we only trained 10M timesteps due to the limitation of computing resources, our method has a higher starting point than the baseline model, and it has been rising. However, the gain of the baselines at the same timestep is not obvious in most cases. This significant improvement is mainly due to two aspects. On the one hand, the adequate knowledge representation facilitates the inference capacity of our model and provides a high-level starting point for warming up the training. On the other hand, the learned subtask planning strategy can make rational decisions while confronting various complicated situations.
5.4 Results on Execution Performance
@@
Table 5: Execution performance on various complicated tasks. We report the average reward with standard deviations of 100 testing episodes with best values in bold.
Ours CORRO PPO SAC SGT DGRL LEAP
GoToSeq 0.30(0.07) 0.23(0.06) 0.14(0.04) 0.05(0.04) 0.15(0.08) 0.19(0.07) 0.17(0.04)
GoToSeqS5R2 0.91(0.02) 0.84(0.02) 0.84(0.03) 0.85(0.02) 0.91(0.04) 0.67(0.19) 0.73(0.06)
SynthSeq 0.22(0.05) 0.16(0.05) 0.12(0.05) 0.12(0.04) 0.18(0.05) 0.15(0.06) 0.14(0.03)
BossLevel 0.26(0.05) 0.15(0.04) 0.14(0.05) 0.12(0.04) 0.19(0.06) 0.13(0.05) 0.18(0.04)
BossLevelNoUnlock 0.24(0.04) 0.17(0.04) 0.10(0.05) 0.04(0.03) 0.11(0.05) 0.12(0.05) 0.16(0.03)
MiniBossLevel 0.29(0.06) 0.26(0.05) 0.25(0.05) 0.06(0.03) 0.14(0.06) 0.22(0.07) 0.23(0.03)
5.5 Results on Various Tree Widths
After evaluating the training and execution performance of our technique, we performed sensitivity analyses on the width of the planning tree. The results are shown in Fig. 8, where we kept the tree depth at $3$, as in the main experiments.
@@ 5.5 Results on Various Tree Widths
(b) BossLevelNoUnlock
(c) MiniBossLevel
(d) GotoSeq
(e) GoToSeqS5R2
(f) SynthSeq
Figure 8: Sensitivity analyses on various tree widths.
The horizontal axis represents the training steps, and the vertical axis shows the average episode reward with a standard deviation, averaged across 10 independent running seeds. Our intuition suggests that a wider tree would result in a more diverse subtask distribution. However, we found that the gain is not the largest when the treewidth is the largest. We speculate that a larger tree width provides more information but weakens the pertinence of problem-solving, leading to confusion. Hence, it is critical to select appropriate treewidth. In the future, we plan to conduct an adaptive selection of the optimal tree width.
5.6 Results on Various Tree Depths
@@
(b) BossLevelNoUnlock
(c) MiniBossLevel
(d) GotoSeq
(e) GoToSeqS5R2
(f) SynthSeq
Figure 9: Sensitivity analyses on various tree depths.
In Fig. 9, we present additional sensitivity analyses to investigate the impact of various tree depths on the behavior of our planning tree. The treewidth is fixed as $2$, the same as the main experiments. The depth represents the model’s capacity to predict the future of the planning tree. Our results show that when the depth is set to $N=2$ or $N=3$, our model achieves comparable performance to the main experiments. However, the curves for $N=4$ show degraded performance compared to the other depths. We speculate that the accumulation of prediction errors about the future at each step can lead the model to make suboptimal decisions, despite the improved foresight ability in making global decisions. Hence, selecting an appropriate depth for the planning tree is crucial for our model’s success, and it is a potential future direction to explore adaptive selection techniques.
5.7 Results on Subtask Execution Statistics
In Fig. 10, we present the subtask proportions on six complex tasks after 1000 episodes of testing. The observed subtask distribution is highly consistent with the predefined task missions, validating the rationality of our approach. Our model focuses on executing the necessary sequences related to the “Go To” task in GoToSeq and GoToSeqS5R2 for most cases. However, occasionally, other subtasks are selected, which we attribute to the inaccuracies of forward prediction while expanding the tree depth. For the remaining four comprehensive tasks, our model executes subtasks at roughly equal frequencies.
Figure 10: The bars represent the mean proportion of each subtask executed by our trained model over 1000 episodes.
6 Conclusion
In this paper, we have considered the difficulties of task planning in solving complicated problems, and propose a two-stage planning framework. The multiple-encoder and individual-predictor regime can extract discriminative subtask representations with dynamic knowledge from priors. The top-k subtask planning tree, which customizes subtask execution plans to guide decisions, is important in complicated tasks. Generalizing our model to unseen and complicated tasks demonstrates significant performance and exhibits greater potential to deploy it to real-life complicated tasks. Overall, our approach provides a promising solution to the task planning problem in complicated scenarios. Future research directions could include automatic learning of explainable subtasks and incorporating global planning as a reference for human decision-making in real-life applications.
Acknowledgement
This work was supported by the National Key R&D Program of China (No.2022ZD0116405), the Strategic Priority Research Program of the Chinese Academy of Sciences (No.XDA27030300), and the National Natural Science Foundation of China under Grant 62073324.
Appendix A Detailed Description about the Intuition for d-UCB
Figure 11: Two instance trees are visualized to verify the effect of d-UCB described in Section 4.4. To ignore the effect of the UCB value, we set the value of all nodes as $1$.
@@ Appendix A Detailed Description about the Intuition for d-UCB
As shown in Fig. 11(a), MS will select path $1$, which obviously places a greater weight on the future with greater uncertainty. However, our d-UCB will choose path $2$, which tends to maximize the immediate benefit by discounting the future uncertain score, which is consistent with our intuition.
Moreover, as shown in Fig.  11(b), unlike MS tending to pick the edge with the highest probability, our d-UCB will choose path $3$. Our d-UCB also considers the optimal subtask to perform in the future when the nearest paths exhibit similar transition probabilities.
Appendix B Detailed Description about Complicated Tasks
(a) ‘BossLevel-v0’
(b) ‘BossLevelNoUnlock-v0’
(c) ‘MiniBossLevel-v0’
(d) ‘GoToSeq-v0’
(e) ‘GoToSeqS5R2-v0’
(f) ‘SynthSeq-v0’
Figure 12: The diagrams for six complicated tasks in BabyAI.
BossLevel.The level requires the agent to implement any union of instructions produced according to Baby Language grammar, it involves the content of all other levels. For example, “open a door and pick up the green key after you go to a ball and put a red box next to the red ball” is a command of the level. This level requires all the competencies including navigating, unblocking, unlocking, opening, going to, picking up, putting next, and so on.
BossLevelNoUnlock.The doors of all rooms are unlocked which the agent can go through without unlocking by a corresponding key. For example, the command “open a purple door, then go to the yellow box and put a yellow key next to a box” does not need the agent to get a key before opening a door. This level requires all the competencies such as BossLevel except for unlocking.
MiniBossLevel.The level is a mini-environment of BossLevel with fewer rooms and a smaller size. This level requires all the competencies such as BossLevel except unlocking.
GoToSeq.The agent needs to complete a series of go-to-object commands in sequence. For example, the command ”go to a grey key and go to a purple key, then go to the green key” only includes go-to instructions. GotoSeq only requires navigating a room or maze, recognizing the location, and understanding the sequence subtasks.
GoToSeqS5R2. The missions are just similar to GoToSeq, which requires the only competency of “go to”. Here, S5 and R2 refer to the size and the rows of the room, respectively.
SynthSeq.The commands, which combine different kinds of instructions, must be executed in order. For example, ”Open the purple door and pick up the blue ball, then put the grey box next to a ball”. This level requires competencies such as GotoSeq except inferring whether to unlock or not.
Appendix C Network Architecture
There is a summary of all the neural networks used in our framework regarding the network structure, layers, and activation functions.
[…]
Table 6: Summary of the Network Architecture
@@ Appendix C Network Architecture
33footnotetext: Please note that ’-’ denotes that we refer readers to check the open source repository about FiLMedBlock and CausalSelfAttention, see https://github.com/caffeinism/FiLM-pytorch and https://github.com/sachiel321/Efficient-Spatio-Temporal-Transformer, respectively.
Jingqing Ruan  received a B.Sc. degree in software engineering from North China Electric Power University, China, in 2019. She is currently a Ph.D. candidate in pattern recognition and intelligent systems at the Institute of Automation, Chinese Academy of Sciences and the University of Chinese Academy of Sciences, China.
Her research focuses on multi-agent reinforcement learning, multi-agent planning on game AI, and graph-based multi-agent coordination.
E-mail: ruanjingqing2019@ia.ac.cn
ORCID ID: $0$-0002-4857-9053
Kaishen Wang  received a B.Sc. degree in electrical engineering and automation from China University of Mining and Technology, China, in 2015. He is currently an M.S. candidate in computer science and technology at the Institute of Automation, Chinese Academy of Sciences and the University of Chinese Academy of Sciences, China.
His research interests include hierarchical reinforcement learning and multi-agent reinforcement learning.
E-mail: wangkaishen2021@ia.ac.cn
ORCID ID: $0$-0002-2787-5873
Qingyang Zhang  received a B.Sc. degree in communication engineering from Shandong University, China, in 2019. She is currently a Ph.D. candidate in pattern recognition and intelligent systems at the Institute of Automation, Chinese Academy of Sciences and the University of Chinese Academy of Sciences, China.
Her research interests include hierarchical reinforcement learning, contrastive learning, representation learning, and multi-agent reinforcement learning.
E-mail: zhangqingyang2019@ia.ac.cn
ORCID ID: $0$-0001-5387-9942
Dengpeng Xing obtained his B.Sc. degree in Mechanical Electronics and M.Sc. degree in Mechanical Manufacturing and Automation from Tianjin University, China, in 2002 and 2006, respectively. He earned his Ph.D. degree in Control Science and Engineering from Shanghai Jiao Tong University, China, in 2010. Currently, he holds the position of Professor at the National Key Laboratory for Multi-modal Artificial Intelligence Systems, Institute of Automation, Chinese Academy of Sciences, China.
Dr. Xing’s research focuses on robot control, reinforcement learning, and causal learning.
E-mail: dengpeng.xing@ia.ac.cn (Corresponding author)
ORCID ID: $0$-0002-8251-9118
Bo Xu  is a professor, the director of the Institute of Automation Chinese Academy of Sciences, and deputy director of the Center for Excellence in Brain Science and Intelligence Technology, Chinese Academy of Sciences.
His main research interests include brain-inspired intelligence, brain-inspired cognitive models, natural language processing and understanding, and brain-inspired robotics.
E-mail: xubo@ia.ac.cn (Corresponding author)
ORCID ID: $0$-0002-$1$-1529